kekeman:

A triangle has sides that measure 7 units, 6 units, and 9.22 units. What is the area of a circle with a circumference that equals the perimeter of the triangle? Use 3.14 for π, and round your answer to the nearest whole number. 22 units2 157 units2 39 units2 70 units2

AZ:

@kekeman What is the perimeter of the triangle? To find the perimeter you just have to add up all the sides of the triangle So what is 7 + 6 + 9.22 = ?

AZ:

The question tells us that the perimeter of this triangle is the same as the circumference of a circle. The formula for circumference of a circle is \( 2\pi r\) The question wants us to find the area of that circle Once you find the perimeter of the triangle (which is equal to the circumference of the circle), we have to then find the radius of the circle and then we can calculate the area from that area of circle = \(\pi r^2\)

kekeman:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ @kekeman What is the perimeter of the triangle? To find the perimeter you just have to add up all the sides of the triangle So what is 7 + 6 + 9.22 = ? \(\color{#0cbb34}{\text{End of Quote}}\) It is 22.22

AZ:

Good So that's also the circumference of the circle and now we need to find the radius of the circle so that we can find the area (which is what the question is asking for) \( 22.22 = 2 \pi r\) can you solve for r? Just divide by \(2 \pi\) on both sides

kekeman:

r = 3.53642283

kekeman:

Is that wrong?

kekeman:

Yeah im not good at math soooo

AZ:

No, that's correct!! Now we just have to find the area of the circle Area = \(\pi r^2\) So basically, plug in r = 3.54 so pi * 3.54^2 =

kekeman:

Idk

AZ:

that ^2 just means you're multiplying that number by itself twice and pi or \(\pi\) is 3.14 so essentially 3.14 * 3.54 * 3.54 = ??

kekeman:

Oh yes oh yes pi is 3.14

dizzymustang156:

what was the warning about

kekeman:

39.349224

AZ:

Good, good Now can you round that to the nearest whole number?

dizzymustang156:

oh right the direct answer sorry

kekeman:

So it would be 39 units2?

dizzymustang156:

yes

AZ:

You got it!

dizzymustang156:

Good job! Good job AZ

kekeman:

Thanks yall i wish i could give the best response to both of yall

AZ:

You're welcome!

snowflake0531:

why would dizzymustag deserve a medal

snowflake0531:

he did nothing but spam

dizzymustang156:

Im a girl keke

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 he did nothing but spam \(\color{#0cbb34}{\text{End of Quote}}\) *She

kekeman:

I am soo sorry

dizzymustang156:

its ok everybody thinks im a boy from my name

dizzymustang156:

and snowflake i don't even have a medal

snowflake0531:

noticed

kekeman:

Now you do :)

dizzymustang156:

thx!

snowflake0531:

Youdo know that hte medal is for 'best answer' she didn't give the best answer, why should she get a medal

kekeman:

Yww

kekeman:

I'm sorry AZ but you really helped a lot

AZ:

sadge

kekeman:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kekeman I'm sorry AZ but you really helped a lot \(\color{#0cbb34}{\text{End of Quote}}\) *sadge

dizzymustang156:

How did AZ get 6 medals?

kekeman:

Lol idk